OpenStudy (anonymous):

(Q2) What will be the effective population size in a panmictic population of 240 with 200 females and 40 polygamous males? Please explain. (A)160 (B)133 (c) 63 (D) 67

OpenStudy (anonymous):

from Wikipedia: When the sex ratio of a population varies from the Fisherian 1:1 ratio, effective population size is given by: N_e^{(v)} = N_e^{(F)} = {4 N_m N_f \over N_m + N_f} Where Nm is the number of males and Nf the number of females. For example, with 80 males and 20 females (an absolute population size of 100): N_e = {4 \times 80 \times 20 \over 80 + 20} ={6400 \over 100} = 64 Again, this results in Ne being less than N.